# On the Average Resistance of $n$-circuits

### v1

@@ Introduction
Figure 2: The first few values of $M_{n}$ connected by a Bézier curves. The dashed line is $y=1$.
@@
$\mathcal{D}(\gamma)$ :-The depth of $\coloneq\ \ \ \ \text{The depth of }\gamma:$
@@ 2 Inverses
$GM\{r(\gamma):\gamma\in\Gamma_{n}\})=\left(\prod_{\gamma\in\widetilde{\Gamma}_{n}}r(\gamma)r(\gamma^{*})\right)^{\frac{1}{Q_{n}}}=\left(\prod_{\gamma\in\widetilde{\Gamma}_{n}}1\right)^{\frac{1}{Q_{n}}}=1$
@@ 4 The Generating Function of $Q_{n}$
The fundamental difference in the nature of the relationships found between $C_{i}(n)$ and $Q_{n}$ in the previous section, allows for a non-trivial recursive expression of $Q_{n}$ in terms of itself. By 3.1.2 and 3.2:
@@
$\mathcal{Q}(x)=\prod\limits_{n=1}^{\infty}{(1-x^{n})^{-q_{n}}}.$ (4)
@@
This result can also be derived using numerical extrapolation on (2), but lots of data is required — analyzing the first $2500$ terms gives a result of $d\approx 3.559$.
@@ 5 Resistance: Circuits & Biscuits
$\begin{split}\widetilde{\varphi}(\frac{a}{b})&=\frac{a}{b}+1=\frac{a+b}{b}\\ \overline{\overline{\varphi}}(\frac{a}{b})&=\frac{1}{\frac{1}{\frac{a}{b}}+\frac{1}{1}}=\frac{a}{a+b}\end{split}$
@@
¯¯$\begin{split}\overline{\overline{\mathfrak{R}}}_{n}&=\sum_{r\in\widetilde{\Psi}_{n-1}}{\overline{\overline{\varphi}}(r)}+\sum_{r\in\overline{\overline{\Psi}}_{n-1}}{\overline{\overline{\varphi}}(r)}\\ &=\sum_{r\in\widetilde{\Psi}_{n-1}}{\overline{\overline{\varphi}}(\frac{r}{1})+\overline{\overline{\varphi}}(\frac{1}{r})}\\ &=\sum_{r\in\widetilde{\Psi}_{n-1}}{\frac{r}{r+1}+\frac{1}{r+1}}\\ &=\mathfrak{q}_{n-1}\\ \end{split}$
@@
$\begin{split}\frac{m}{\sum_{i=1}^{m}{\frac{1}{a_{i}+b_{j}}}}&\leq\frac{1}{m}\sum_{i=1}^{m}{a_{i}+b_{j}}\\ \Rightarrow\frac{m}{\frac{1}{m}\sum_{i=1}^{m}{a_{i}+b_{j}}}&\leq\sum_{i=1}^{m}{\frac{1}{a_{i}+b_{j}}}\\ \Rightarrow m-\sum_{i=1}^{m}{\frac{b_{j}}{a_{i}+b_{j}}}&\leq m-\frac{b_{j}m}{\frac{1}{m}\sum_{i=1}^{m}{a_{i}+b_{j}}}\\ \Rightarrow\sum_{i=1}^{m}{\left(1-\frac{b_{j}}{a_{i}+b_{j}}\right)}&\leq\frac{\sum_{i=1}^{m}{(a_{i}+b_{j})}-b_{j}m}{\frac{1}{m}\sum_{i=1}^{m}{a_{i}+b_{j}}}\\ \Rightarrow\sum_{i=1}^{m}{\frac{a_{i}}{a_{i}+b_{j}}}&\leq\frac{\sum_{i=1}^{m}{a_{i}}}{\frac{1}{m}\sum_{i=1}^{m}{a_{i}+b_{j}}}\\ \Rightarrow\frac{1}{m}\sum_{i=1}^{m}{\frac{a_{i}b_{j}}{a_{i}+b_{j}}}&\leq\frac{\frac{1}{m}\sum_{i=1}^{m}{a_{i}b_{j}}}{\sum_{i=1}^{m}{a_{i}+b_{j}}}\\ \Rightarrow\frac{1}{m}\sum_{i=1}^{m}{{\frac{1}{\frac{1}{a_{i}}+\frac{1}{b_{j}}}}}&\leq\frac{1}{\frac{1}{\frac{1}{m}\sum_{i=1}^{m}{a_{i}}}+\frac{1}{b_{j}}}\end{split}$
@@
Begin by noting that any $n$-biscuit has at most one non-unit sub-circuit; this is because, by definition, the construction of an $n$-biscuit is invertible; hence, if one assumes, by way of contradiction that there exists an $n$-biscuit with more than one non-unit sub-circuit, after deconstructing all of its unit sub-circuits, one can remove neither a series nor a parallel connection to a unit circuit, and therefore, reaches a contradiction.
Now consider a parallel $n$-biscuit $\beta$ with $k$ unit sub-circuits and a non-unit sub-circuit that is a series $(n-k)$-biscuit. As by 5.1 the connection of the $k$ unit circuits in parallel is the smallest possible value for the resistance of a $k$-circuit, any first-depth rearrangement of this biscuit will result in a positive change in its total resistance. Inversely, any such rearrangement in a series biscuit will result in negative change; this corresponds to a second-depth rearrangement of $\beta$. The argument can be recursively repeated to show that any odd-depth rearrangement of parallel $n$-biscuits results in a positive change in total resistance and any even-depth rearrangement in a negative; the reverse is true for series circuits. It can be shown, using calculus, that the magnitude of change decreases as depth increases, resulting in the statement of the theorem. The rigorous proof is left as an exercise for the reader. ∎
@@
Any parallel $n$-circuit can be thought of as a rearrangement of a parallel $n$-biscuit, and 5.5 shows that the average resistance of these rearrangements is higher than that of the parallel $n$-biscuits, resulting in the lower bound.
@@
$\begin{split}\overline{\overline{M}}_{n}&\leq\frac{1}{\sum_{k=1}^{n-1}{Q_{k}\mathfrak{q}_{n-k}}}\sum_{k=1}^{n-1}\frac{Q_{k}\mathfrak{q}_{n-k}M_{k}\widetilde{\mathfrak{M}}_{n-k}}{M_{k}+\widetilde{\mathfrak{M}}_{n-k}}\\ &=\frac{1}{\sum_{k=1}^{n-1}{Q_{k}2^{-k}}}\sum_{k=1}^{n-1}\frac{2^{-k}Q_{k}M_{k}\widetilde{\mathfrak{M}}_{n-k}}{M_{k}+\widetilde{\mathfrak{M}}_{n-k}}\\ &<\frac{\sum_{k=1}^{n-1}\frac{5M_{k}}{2M_{k}+5}Q_{k}2^{-k}}{\sum_{k=1}^{n-1}{Q_{k}2^{-k}}}\leq\max{\{\frac{5M_{k}}{2M_{k}+5}\}}<\frac{5}{2}\end{split}$
@@ 6.2 More $C_{i}(n)$ Properties
Although peculiar to interpret its meaning,

### v2

@@ Introduction
Figure 2: The first few values of $M_{n}$ connected by Bézier curves. The dashed line is $y=1$.
@@
$\mathcal{D}(\gamma)$ :-The depth of the circuit $\coloneq\ \ \ \ \text{The depth of the circuit }\gamma:$
@@ 2 Inverses
$GM(\{r(\gamma):\gamma\in\Gamma_{n}\})=\left(\prod_{\gamma\in\widetilde{\Gamma}_{n}}r(\gamma)r(\gamma^{*})\right)^{\frac{1}{Q_{n}}}=\left(\prod_{\gamma\in\widetilde{\Gamma}_{n}}1\right)^{\frac{1}{Q_{n}}}=1$
@@ 4 The Generating Function of $Q_{n}$
The fundamental difference in the nature of the relationships found between $C_{i}(n)$ and $Q_{n}$ in the previous section allows for a non-trivial recursive expression of $Q_{n}$. By 3.1.2 and 3.2:
@@
$\mathcal{Q}(x)=\prod\limits_{n=1}^{\infty}{(1-x^{n})^{-q_{n}}}$ (4)
@@
This result can also be derived using numerical extrapolation on (2), but a lot of data is required — analyzing the first $2500$ terms gives a result of $d\approx 3.559$.
@@ 5 Resistance: Circuits & Biscuits
$\begin{split}\widetilde{\varphi}\left(\frac{a}{b}\right)&=\frac{a}{b}+1=\frac{a+b}{b}\\ \overline{\overline{\varphi}}\left(\frac{a}{b}\right)&=\frac{1}{\frac{1}{\frac{a}{b}}+\frac{1}{1}}=\frac{a}{a+b}\end{split}$
@@
¯¯$\begin{split}\overline{\overline{\mathfrak{R}}}_{n}&=\sum_{r\in\widetilde{\Psi}_{n-1}}{\overline{\overline{\varphi}}(r)}+\sum_{r\in\overline{\overline{\Psi}}_{n-1}}{\overline{\overline{\varphi}}(r)}\\ &=\sum_{r\in\widetilde{\Psi}_{n-1}}{\overline{\overline{\varphi}}\left(\frac{r}{1}\right)+\overline{\overline{\varphi}}\left(\frac{1}{r}\right)}\\ &=\sum_{r\in\widetilde{\Psi}_{n-1}}{\frac{r}{r+1}+\frac{1}{r+1}}\\ &=\mathfrak{q}_{n-1}\\ \end{split}$
@@
$\begin{split}\frac{m}{\sum_{i=1}^{m}{\frac{1}{a_{i}+b_{j}}}}&\leq\frac{1}{m}\sum_{i=1}^{m}{a_{i}+b_{j}}\\ \Rightarrow\frac{m}{\frac{1}{m}\sum_{i=1}^{m}{a_{i}+b_{j}}}&\leq\sum_{i=1}^{m}{\frac{1}{a_{i}+b_{j}}}\\ \Rightarrow m-\sum_{i=1}^{m}{\frac{b_{j}}{a_{i}+b_{j}}}&\leq m-\frac{b_{j}m}{\frac{1}{m}\sum_{i=1}^{m}{a_{i}+b_{j}}}\\ \Rightarrow\sum_{i=1}^{m}{\left(1-\frac{b_{j}}{a_{i}+b_{j}}\right)}&\leq\frac{\sum_{i=1}^{m}{(a_{i}+b_{j})}-b_{j}m}{\frac{1}{m}\sum_{i=1}^{m}{a_{i}+b_{j}}}\\ \Rightarrow\sum_{i=1}^{m}{\frac{a_{i}}{a_{i}+b_{j}}}&\leq\frac{\sum_{i=1}^{m}{a_{i}}}{\frac{1}{m}\sum_{i=1}^{m}{a_{i}+b_{j}}}\\ \Rightarrow\frac{1}{m}\sum_{i=1}^{m}{\frac{a_{i}b_{j}}{a_{i}+b_{j}}}&\leq\frac{\frac{1}{m}\sum_{i=1}^{m}{a_{i}b_{j}}}{\frac{1}{m}\left(\sum_{i=1}^{m}{a_{i}}\right)+b_{j}}\\ \Rightarrow\frac{1}{m}\sum_{i=1}^{m}{{\frac{1}{\frac{1}{a_{i}}+\frac{1}{b_{j}}}}}&\leq\frac{1}{\frac{1}{\frac{1}{m}\sum_{i=1}^{m}{a_{i}}}+\frac{1}{b_{j}}}\end{split}$
@@
Begin by noting that any $n$-biscuit has at most one non-unit sub-circuit; this is because, by definition, the construction of an $n$-biscuit is invertible. Hence, if one assumes by way of contradiction that there exists an $n$-biscuit with more than one non-unit sub-circuit, after reverting the construction of all of its unit sub-circuits, one can remove neither a series nor a parallel connection to a unit circuit, reaching a contradiction.
Now consider a parallel $n$-biscuit $\beta$ with $k$ unit sub-circuits and a non-unit sub-circuit that is a series $(n-k)$-biscuit. As by 5.1 the connection of the $k$ unit circuits in parallel is the smallest possible value for the resistance of a $k$-circuit, any first-depth rearrangement of this biscuit will result in a positive change in its total resistance. Inversely, any such rearrangement in a series biscuit will result in a negative change; this corresponds to a second-depth rearrangement of $\beta$. This argument can be recursively repeated to show that any odd-depth rearrangement of parallel $n$-biscuits results in a positive change in total resistance and any even-depth rearrangement in a negative; the converse is true for series circuits. It can be shown, using calculus, that the magnitude of change decreases as depth increases, resulting in the statement of the theorem. The rigorous proof is left as an exercise for the reader. ∎
@@
Any parallel $n$-circuit can be thought of as a rearrangement of a parallel $n$-biscuit, and 5.5 shows that the average resistance of these rearrangements is higher than that of parallel $n$-biscuits, resulting in the lower bound.
@@
$\begin{split}\overline{\overline{M}}_{n}&\leq\frac{1}{\sum_{k=1}^{n-1}{Q_{k}\mathfrak{q}_{n-k}}}\sum_{k=1}^{n-1}\frac{Q_{k}\mathfrak{q}_{n-k}M_{k}\widetilde{\mathfrak{M}}_{n-k}}{M_{k}+\widetilde{\mathfrak{M}}_{n-k}}\\ &=\frac{1}{\sum_{k=1}^{n-1}{Q_{k}2^{-k}}}\sum_{k=1}^{n-1}\frac{2^{-k}Q_{k}M_{k}\widetilde{\mathfrak{M}}_{n-k}}{M_{k}+\widetilde{\mathfrak{M}}_{n-k}}\\ &<\frac{\sum_{k=1}^{n-1}\frac{5M_{k}}{2M_{k}+5}Q_{k}2^{-k}}{\sum_{k=1}^{n-1}{Q_{k}2^{-k}}}\leq\max{\left\{\frac{5M_{k}}{2M_{k}+5}\right\}}<\frac{5}{2}\end{split}$
@@ 6.2 More $C_{i}(n)$ Properties
Although difficult to interpret its meaning,